OpenStudy (anonymous):

For question 4 and 5, what values for Ɵ (0≤Ɵ≤2π) satisfy the equation? 4. 4cosƟ+1= 2cosƟ (1 point) a. 2π/3, 4π/3 b. π/6, 5π/6 c. 7π/6, 11π/6 d. π/3, 6π/3 5. cosƟ- tanƟ cosƟ=0 (1 point) a. 0, π/4, π, 5π/4 b. π/4, 5π/4 c. π/2, 3π/4, 3π/2, 7π/4 d. π/2, 7π/6, 3π/2, 11π/6

OpenStudy (anonymous):

I really need help with this, I do not understand this. If you can please show me step-by-step how you did it I would be happy. Thank you in advance.

OpenStudy (anonymous):

Ok, for question four you have that: 4cos(theta)+1=2cos(theta) 2cos(theta)+1=0 2cos(theta)=-1 cos(theta)=-1/2 And for question five: cos(theta)-tan(theta)cos(theta)=0 If you divide both sides for cos(theta) you will have 1-tan(theta)=0 tan(theta)=1 Do you know how to get the values of the angles now?

OpenStudy (anonymous):

Not really.. I'm sorry, I really don't understand this stuff.

OpenStudy (anonymous):

Did you understand the steps until now?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Do I use the unit circle?

OpenStudy (anonymous):

Do you know what angle makes cos(theta) = 1/2?

OpenStudy (anonymous):

No. I am really bad at math.

OpenStudy (anonymous):

I'm sorry, but I really do need your help.

OpenStudy (anonymous):

Ok, if I where you I will learn the values for some angles from 0 to 90º Here is a table: http://marcsmnm.blogspot.com.es/2011/03/memorizing-common-trig-values.html From these values you can compute the others. You know that cosine is postive in the first and the fourth quadrant of a circle, and negative in the second and the third ok?

OpenStudy (anonymous):

Is the answer B? I was looking at the Unit Circle and I thought that might be right.

OpenStudy (anonymous):

Now, if we know that for cos(theta)=1/2 only if theta=30º=pi/6, the values when its negative are pi/6+pi/2 and pi/6+pi so for question 4 answer is a). Is that ok? Can you do it for question 5?

OpenStudy (anonymous):

Ummm. Okay, I'm not sure.

OpenStudy (anonymous):

I turned it in and you where right on #4 it was a. #5 was B

OpenStudy (anonymous):

Yes, it is. Good job!